OpenStudy (gamer56):

What is the value of the expression when a = 6, b = 4, and c = 8? 2a/3b-c

OpenStudy (joannablackwelder):

Is this \[\frac{ 2a }{ 3b-c }\] or \[\frac{ 2a }{ 3a }-c\] ?

OpenStudy (joannablackwelder):

@gamer56 ?

OpenStudy (joannablackwelder):

Sorry, the second one should be \[\frac{ 2a }{ 3b }-c\]

OpenStudy (gamer56):

its the first one

OpenStudy (gamer56):

@JoannaBlackwelder

OpenStudy (joannablackwelder):

Ok, cool. Substitute your values for a, b, and c into the expression and simplify.

OpenStudy (joannablackwelder):

\[\frac{ 2*(6) }{ 3*(4)-8 }\]

OpenStudy (joannablackwelder):

Make sense so far?

OpenStudy (joannablackwelder):

Well, use the order of operations to simplify from here. Let me know what you get. :)

OpenStudy (gamer56):

@KlOwNlOvE can you check my answer? i think its 4

OpenStudy (joannablackwelder):

Close, but not quite. \[\frac{ 12 }{ 12-8 }\]

OpenStudy (joannablackwelder):

\[\frac{ 12 }{ 4 }=?\]